# Revisiting In-context Learning Inference Circuit in Large Language Models

### v3

@@ 5.3 Explaination towards Observed ICL Phenomena
Prediction Bias. (1) Contextual Bias: As shown in §5.2 and Appendix E, direct decoding insufficiently encoded information adds meanless logits into LM’s output, causing a background prediction value even if no queries are given (named bias). (2) Positional Bias: As shown in §3.2, closer input texts are encoded more similarly, so label tokens near the query have more similar information to the query, causing more attention assignment in the induction processing, so that more influences on the prediction. (3) Frequency Bias: As shown in §4.2, in the induction, some attention heads are without selectivity towards labels, causing an averaged copy processing from label tokens to the query, triggering a prediction bias towards the label frequency in the demonstration, even if their contribution (absolute value of attention score on label tokens) is small. All three biases are observed by Zhao et al. (2021), and can be removed by ICL calibration methods.
@@ 6 Conclusion and Discussion
Conclusion. In summary, this paper restores ICL inference into 3 basic operations and confirms their existence. Careful measurements are conducted to capture and explain various phenomena successfully. Moreover, ablation studies show the proposed inference circuit dominates and reveals the existence of bypass mechanisms. We hope this paper can bring new insight into ICL practice.
@@
Inference Acc. # Para. Speed
@@
Full + Cent. 83.24% 69.5B 1.00$\times$
@@
Layer34 + Cent. 84.27% 31.2B 2.38$\times$
The Role of Early and Later Layers. Our framework and Fig. 3 show: encoding result of Step 1 can be directly used for classification with reliable decoding, and later transformer layers are not contributing to centroid classification accuracies, leading to a taxonomy of Encoding for Step 1 and Output Preparation for Step 2 and 3: LMs complete multi-task classification implicitly in early layers, and verbalize it by merging task-specific label semantics in later layers. Therefore, we suggest an early-exiting inference: removing some top layers and using a centroid classifier (Cho et al., 2024) to accelerate ICL as shown in Table 3 and Appendix E.
@@
Pre-training Possibility from Natural Language Data. A large gap can be considered between such a delicate circuit and gradient descent pre-training on the wild data. However, we believe the wild training target contains the ICL circuit functionally. Based on the previous works finding trainability of ICL on linear representation-label pairs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b), we speculate that in early training step, Transformers learn to extract linear representations shown in §3 from wild data (Appendix B), serving as the training input of later layers to evoke the emergence of induction heads with the same mechanism shown in aforementioned previous works. Moreover, our conclusion of Step 3 highlights the input data requirements for the later layers: these data should activate the multiplex of hidden space, i.e., it should implicate multi-task classification with a wide distribution, which is consistent with the aforementioned previous works.
@@
Comparison with Previous Works. Several prior studies have sought to interpret ICL as known algorithms, including implicit gradient descent (Dai et al., 2023), kernel regression (Han et al., 2023a), and implicit Bayesian inference (Xie et al., 2022), etc. However, as noted in §2.1, these approaches fall short of fully explaining the phenomena observed during the ICL inference process. For instance, gradient descent is known to be fragile against label noise (Zhang et al., 2021), leading to a misalignment when analogies are drawn to ICL, which is robust against label noise (Min et al., 2022). Similarly, attempts to explain ICL as kernel regression fail to account for positional bias, making them disconnected from empirical studies that document various inference phenomena in real-world LMs. Also, other works have employed induction circuits to explain ICL dynamics (Wang et al., 2023; Elhage et al., 2021; Olsson et al., 2022), yet significant gaps remain in aligning these explanations with empirical observations. Our work is the first to unify the fragmented conclusions from prior empirical studies (discussed in §5.3) through detailed experimental measurements. By demonstrating the alignment of these observations, we emphasize the novelty and primary contribution of our paper.
@@
Limitations. (1) These 3 basic operations are not functionally indivisible. Ideally, one can reduce every operation in ICL inference to the interconnection of special attention heads to ulteriorly examine how the operating subspaces interact between steps, and reconstruct ICL behavior from a minimal set of attention heads. Also, although we show the significance of the inference circuit in the ablation analysis (§5.1), measuring the connectivity of these attention heads can also be beneficial to get more insights into the circuit. (2) The conclusions may not align with scenarios where ground-truth labels are not provided in the context in some cases, which are often referred to as in-weight learning, where significant differences or even antagonism with standard ICL have been highlighted by previous works (Chan et al., 2022; Reddy, 2024) [^9] In these works, toy models even struggle to work on both inputs simultaneously, so that it is natural to consider different inference circuit for the in-weight learning., reasonably and necessarily warranting separate discussion (discussed in Appendix G). This paper explains the inference behavior of the model under ICL conditions, leaving the in-weight learning scenario for future works. (3) We only focus on classification tasks, while we believe that our findings can be applied to non-classification tasks, efforts are still needed to fill the gap.
Reproducibility Statement
The official code implementation of this paper by the author can be found in https://github.com/hc495/ICL_Circuit.
@@
N.I., as the supervisor, provided valuable feedback, revisions, and financial support in research. M.K. and Y.S. made minor contributions that placed them on the borderline of authorship criteria: they participated in discussions and offered comments, some of which were incorporated.
@@ Acknowledgments
We sincerely thank the chairs and reviewers of ICLR 2025 for their thoughtful and insightful feedback on this paper, their perceptive reviews and discussions have greatly improved this work.
@@ A.1 Detailed Overall Experimental Settings
Other. All the experiment materials (models and datasets) are loaded from huggingface. For the BGE M3, we use its pooler_output as the output feature.
@@ A.2 Calculation of Mutual Nearest-neighbor Kernel Alignment
In this paper, we need to measure the similarity between features from two different models or model layers. There are many approaches (Klabunde et al., 2023), and we use mutual nearest-neighbor kernel alignment (Huh et al., 2024), which is relatively efficient and accurate, calculated as follows to measure the similarity of representation from the same object set $\mathcal{X}=\left\{x_{i}\right\}_{i=1}^{n}$ in different feature spaces.
Given a representation mapping $\delta:\mathcal{X}\rightarrow\mathbb{H}^{d}$ from the objects to a space where similarity measurement $\left\langle\cdot,\cdot\right\rangle:\mathbb{H}^{d}\times\mathbb{H}^{d}\rightarrow\mathbb{R}$ is defined, we can calculate the similarity map from dataset $\mathcal{X}$ as $\mathcal{S}_{\delta}\in\mathbb{R}^{n\times n}$, where the elements are $\mathcal{S}_{\delta|i,j}=\langle\delta\left(x_{i}\right),\delta\left(x_{j}\right)\rangle$, especially, we axiomatic define $\left\langle x,x\right\rangle=1$, so we set the diagonal element $\mathcal{S}_{\delta|i,i}\fallingdotseq 0$ since they are trivial values.
Given two encoding $\delta_{1}$ and $\delta_{2}$, two similarity map can be calculated as $\mathcal{S}_{\delta_{1}}$ and $\mathcal{S}_{\delta_{2}}$ on the same object set $\mathcal{X}$. For each line vector index $i=1,2,\dots,n$ in $\mathcal{S}_{\delta_{1}}$, we select the index of top-$k$ elements from greater to lower as $\mathrm{top}_{k}\left(\mathcal{S}_{\delta_{1}|i}\right)$. Similarly, we get $\mathrm{top}_{k}\left(\mathcal{S}_{\delta_{2}|i}\right)$ from $\mathcal{S}_{\delta_{2}}$.
@@
Figure 9: Distributions (clusters) of representations generated by: Left: encoder model (BGE), clustering w.r.t. the label; Middle: ICL on the forerunner token, where representations gather into one point when no Transformer operation is conducted; Right: ICL on the label token, where representations gather into points w.r.t. label (using a 2-way example) when no Transformer operation is conducted, causing a high background value.
@@
Given two specific tokens $x_{i}$ and $x_{j}$ where kernel alignment is calculated from different ICL-styled input sequences $p_{i}$ and $p_{j}$, in a specific layer of a decoder Transformer, the representations can be written as $\delta(x)=e(x)+\epsilon(p)$, where $e(x)$ are the embedding vector of the token $x$, and $\epsilon(p)$ are the residual side-flow w.r.t the context $p$.
Intuition. As shown in Fig. 9, in the hidden states of ICL, the hidden state on the label token has a prior clustering, making it naturally similar to the representation generated by the encoder model, even if the ICL process does not encode it sufficiently. So, at layer 0, since the model is not able to perform any encoding in this layer, the kernel alignment on the forerunner is on a random baseline value of $0.125$, but the value on the label token will be greater. In the case where (1) BGE generates fully linearly separable clusters with sufficient inter-cluster distance, and (2) the number of samples for each label is not less than 64, this upper bias can be expected to be $0.125(|\mathbb{Y}|-1)$ (proof omitted). Intuitively, such bias can propagate along the residual network to every layer, making the results of any layer unfaithful.
@@
Background Values of Kernel Alignment. According to the explanation above, it is intuitive to conclude that $\mathrm{top}_{k}\left(\mathcal{S}_{\delta|i}\right)$ from label tokens is easier to cluster samples with the same label as $y_{i}$. Moreover, a well-pre-trained encoder can catch the prior distribution determined by these labels, and also cluster samples with the same label, causing a high similarity of similarity map, so that a high but unfaithful kernel alignment as the background value from the similarity on $e(y)$ but not $\epsilon(p)$. An intuitive verification of such background value is shown in the Fig. 2 (Left), where the “Label Token” curve has a high value in layer 0 with $\epsilon(p)=0$. However, the background value also indicates that the representation generated by BGE correctly clusters the samples, confirming its reliability.
@@ A.3 Training and Inference of Centroid Classifier
In this paper, we follow Cho et al. (2024) to train centroid classifiers as a probe toward hidden states of LMs. In detail, given the LM’s hidden states set $\left\{h_{i}^{l}\right\}_{i=1}^{m}$ of the selected tokens (according to the experimental setting, the last label token or forerunner token) in layer $l$ from an [input prompt]-[query label] set $\mathcal{Z}=\left\{\left(p_{i},y_{i}\right)\right\}_{i=1}^{m}$, where the labels are limited in label space $\mathbb{Y}$, in the training phase, we calculate the centroid of the hidden state $\bar{h}^{l}_{y}$ for each label respectively:
@@
Implementation. In our experiments, we set training sample number $m\fallingdotseq 256$, similarity function $\left\langle a,b\right\rangle=-\|a-b\|_{2}$.
@@
In Fig. 6 (Middle), we define a Correct Label Assignment, here we introduce how this measurement is calculated. Suppose we have an attention score $\mathcal{A}_{W,f}(K,Q)$ calculated as:
@@
For Fig. 8, we calculate the magnitude of Step 1 as the finite differences of kernel alignment in Fig. 2 (Left, Forerunner Token of Label). We directly use the head counting of Fig. 5 (Middle) and Fig. 6 (Left) as the magnitude of Steps 2 and 3. These data are regularized and converted into transparencies.
Figure 10: Visualization of non-trivial attention connection defined and disconnected in ablation analysis in §5.1 and Table 1. Notations are same as Table 1.
@@ A.7 Experiment Setting of Ablation Experiments in Table 1
For reference, for each experiment, we also conduct controlled experiments where the same amount of randomly selected attention connections are removed from the same layers as the experimental values. The controlled results are shown as smaller numbers under the experimental results.
@@ Appendix B LM Pre-training Dynamics Measured by ICL Circuit
Figure 11: Operating magnitude (normalized) and ICL accuracy w.r.t. pre-training steps on Pythia 6.9B and SST-2.
One can divide a self-regression model into an early part and a later part, where the early part encodes the input into a hidden representation, and the later part decodes the hidden representation back to the input. So, the training object can also be divided into an encoding loss and a decoding loss. According to the discussion in §6, the operation of Step 1 can be classified as encoding, and the other two steps of the induction circuit can be classified as decoding. Intuitively, since the decoding operations require the encoding results as input, unless the encoding operation converges to a stable output, the decoding can not be trained since the input-output mapping is noised, causing unstable gradients to interfere with the training (Liu et al., 2020).
We confirm such inference by a measurement in Pythia 6.9B (Biderman et al., 2023) as shown in Fig 11, where: (1). The magnitude of the 3 operations is emergent in the early phase of pre-training (less than 10k steps), and is monotonically increasing, while the encoding operation has the fastest growth rate. (2). ICL capacity appearance after all three operations reaches a high level (around 50k steps, notice that the random accuracy is 0.5), while the curve morphology of the operating magnitude against the layer number (shown in Fig. 12) is convergence to the last training step and also the main results in this paper. Such results suggest that: LMs start to produce the inner encoding in the very early steps of the pre-training, and can be an important fundamental in building the subsequent induction circuits, as explained in the previous works mentioned in §2.1.

### v4

@@ 5.3 Explaination towards Observed ICL Phenomena
Prediction Bias. (1) Contextual Bias: As mentioned in §5.2 and shown in Appendix E, direct decoding insufficiently encoded information adds meanless logits into LM’s output, causing a background prediction value even if no queries are given (named bias). (2) Positional Bias: As shown in §3.2, closer demonstrations are encoded more similarly, so label tokens near the query have more similar information to the query, causing more attention assignment in the induction processing, so that more influences on the prediction. (3) Frequency Bias: As shown in §4.2, in the induction, some attention heads are without correct selectivity towards labels, causing an averaged copy processing from label tokens to the query, triggering a prediction bias towards the label frequency in the demonstration, even if their contribution (absolute value of attention score on label tokens) is small. All three biases are observed by Zhao et al. (2021), and can be removed by ICL calibration methods.
@@ 6 Conclusion and Discussion
Conclusion. In summary, this paper restores ICL inference into 3 basic operations and confirms their existence. Fine-grained measurements are conducted to capture and explain various phenomena successfully. Moreover, ablation studies show the proposed inference circuit dominates and reveals the existence of bypass mechanisms. We hope this paper can bring new insight into ICL practice.
The Role of Early and Later Layers. Our framework shows: The encoding result of Step 1 can be directly used for classification with reliable decoding, and later transformer layers are not contributing to centroid classification accuracies (Fig. 3), leading to a taxonomy of Input Encoding for Step 1 and Output Preparation for Step 2 and 3: LMs complete multi-task classification implicitly in early layers, and verbalize it by merging task-specific label semantics in later layers. Therefore, we suggest an early-exiting inference: removing some top layers and using a centroid classifier (Cho et al., 2024) to accelerate ICL as shown in Table 3 and Appendix E.
@@
Inference Acc. # Param. Speed
@@
Full + Centroid 83.24% 69.5B 1.00$\times$
@@
Layer34 + Centroid 84.27% 31.2B 2.38$\times$
@@
Pre-training Possibility from Natural Language Data. A large gap can be considered between such a delicate circuit and gradient descent pre-training on the wild data. However, we believe the wild training target contains the ICL circuit functionally: Based on the previous works finding trainability of ICL on simplified linear representation-label pairs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b), we speculate that in early training step, Transformers learn to extract linear representations shown in §3 from wild data (detailed in Appendix B), serving as the training input of later layers to evoke the emergence of induction heads with the same mechanism shown in aforementioned previous works. Moreover, our conclusion of Step 3 highlights the input data requirements for the later layers: These data should activate the multiplex of hidden space, i.e., it should implicate multi-task classification with a wide distribution, which is consistent with the aforementioned previous works.
@@
Comparison with Previous Works. Several prior studies have sought to interpret ICL as known processes, including implicit gradient descent (Dai et al., 2023), kernel regression (Han et al., 2023a), and implicit Bayesian inference (Xie et al., 2022), etc. However, as stated in §2.1, these approaches fall short of fully explaining the phenomena observed during the ICL inference process. For instance, gradient descent is known to be fragile against label noise (Zhang et al., 2021), leading to a misalignment when analogies are drawn to ICL, which is robust against label noise (Min et al., 2022; Liu et al., 2020). Similarly, attempts to explain ICL as kernel regression fail to account for positional bias, making them disconnected from empirical studies that have observed various inference phenomena in real-world LMs. Also, other works have employed induction circuits to explain ICL dynamics (Wang et al., 2023; Elhage et al., 2021; Olsson et al., 2022), yet significant gaps remain in aligning these explanations with empirical observations. Our work is the first to unify the fragmented conclusions from prior empirical studies (discussed in §5.3) through detailed experimental measurements. By demonstrating the alignment of these observations, we emphasize the novelty and primary contribution of our paper.
@@
Limitations & Open-questions. (1) These 3 basic operations are not functionally indivisible. Ideally, mechanistic interpretability aims to reduce every operation in ICL inference to the interconnection of special attention heads to ulteriorly examine how the operating subspaces interact between steps, and reconstruct ICL behavior from a minimal set of attention heads. Also, although we show the significance of the inference circuit in the ablation analysis (§5.1), measuring the connectivity of these attention heads can also be beneficial to get more insights into the circuit. (2) The conclusions may not align with scenarios where ground-truth labels are not provided in the context in some cases, which are often referred to as in-weight learning, where significant differences or even antagonism with standard ICL have been highlighted by previous works [^9] In these works, toy models even struggle to work on both inputs simultaneously, so that it is natural to consider different inference mechanism for the so-called in-weight learning. (Chan et al., 2022; Reddy, 2024), reasonably and necessarily warranting separate discussion (discussed in Appendix G). This paper explains the inference behavior of the model under ICL conditions, leaving the in-weight learning scenario for future works. (3) We only focus on classification tasks, while we believe that our findings can be applied to non-classification tasks, efforts are still needed to fill the gap.
@@
Naoya Inoue, as the supervisor, provided valuable feedback, revisions, and financial support for this paper. M.K. and Y.S. made minor contributions that placed them on the borderline of authorship criteria: they participated in discussions and offered comments, some of which were incorporated.
Reproducibility Statement
[…]
@@ Acknowledgments
The authors would like to thank Mr. Lihao Liu at the Beijing Institute of Technology, and Ms. Yunpin Li at the Capital Normal University for proofreading.
The authors would like to sincerely thank the chairs and reviewers of ICLR 2025 for their thoughtful and insightful feedback on this paper, their perceptive reviews and discussions have greatly improved this work.
@@ A.1 Detailed Overall Experimental Settings
Other. All the experiment materials (models and datasets) are downloaded from huggingface. For the BGE M3, we use its pooler_output as the output feature.
@@ A.2 Calculation of Mutual Nearest-neighbor Kernel Alignment
In this paper, we need to measure the similarity between features from two different models or model layers. There are many approaches (Klabunde et al., 2023), and we use mutual nearest-neighbor kernel alignment (Huh et al., 2024), which is relatively concise and accurate, calculated as follows to measure the similarity of representation from the same object set $\mathcal{X}=\left\{x_{i}\right\}_{i=1}^{n}$ in different feature spaces.
Calculation of Mutual Nearest-neighbor Kernel Alignment. Given a representation mapping $\delta:\mathcal{X}\rightarrow\mathbb{H}^{d}$ from the objects to a space where similarity measurement $\left\langle\cdot,\cdot\right\rangle:\mathbb{H}^{d}\times\mathbb{H}^{d}\rightarrow\mathbb{R}$ is defined, we can calculate the similarity map from dataset $\mathcal{X}$ as $\mathcal{S}_{\delta}\in\mathbb{R}^{n\times n}$, where the elements are $\mathcal{S}_{\delta|i,j}=\langle\delta\left(x_{i}\right),\delta\left(x_{j}\right)\rangle$, especially, we axiomatic define $\left\langle x,x\right\rangle=1$, so we set the diagonal element $\mathcal{S}_{\delta|i,i}\fallingdotseq 0$ since they are trivial values that disturb the following calculation with values $1$.
Given two representation functions $\delta_{1}$ and $\delta_{2}$, two similarity map can be calculated as $\mathcal{S}_{\delta_{1}}$ and $\mathcal{S}_{\delta_{2}}$ on the same object set $\mathcal{X}$. For each line vector index $i=1,2,\dots,n$ in $\mathcal{S}_{\delta_{1}}$, we select the index of top-$K$ elements from greater to lower as $\mathrm{top}_{k}\left(\mathcal{S}_{\delta_{1}|i}\right)$. Similarly, we get $\mathrm{top}_{K}\left(\mathcal{S}_{\delta_{2}|i}\right)$ from $\mathcal{S}_{\delta_{2}}$.
@@
[…]
@@
In a specific layer of a decoder Transformer, the representations of a token $x$ can be written as $\delta(x)=e(x)+\epsilon(p)$, where $e(x)$ is the embedding vector of the token $x$, and $\epsilon(p)$ are the residual side-flow w.r.t. the context $p$. Given two specific tokens $x_{i}$ and $x_{j}$ where kernel alignment is calculated from different ICL-styled input sequences $p_{i}$ and $p_{j}$, we discuss the background value of kernel alignment as below.
Intuition. As shown in Fig. 9, the hidden states on the label token have a prior clustering, making them naturally similar to the representation generated by the encoder model, even if the ICL process does not encode it sufficiently. So, at layer 0, since the model is not able to perform any contextual encoding in this layer, the kernel alignment on the forerunner is on a random baseline value of $0.125$, but the value on the label token will be greater (estimated below). In the case where (1) BGE generates fully linearly separable clusters against $x$ in each class, with sufficient inter-cluster distance, and (2) the number of samples for each label is not less than 64, such upper bias can be expected to be $0.125(|\mathbb{Y}|-1)$ (proof omitted). Intuitively, such bias can propagate along the residual network from the embedding of $[y]$ to every layer, making the results of every layer upper-biased.
@@
Background Values of Kernel Alignment. According to the explanation above, as shown in Fig. 9, it is intuitive to conclude that $\mathrm{top}_{K}\left(\mathcal{S}_{\delta|i}\right)$ from label tokens is easier to cluster samples with the same label as $y_{i}$. Moreover, a well-pre-trained encoder can catch the prior distribution of the input texts determined by their labels, and also cluster samples with the same label, causing a high similarity of similarity map, so that a high but unfaithful kernel alignment as the background value from the similarity on $e(y)$ but not $\epsilon(p)$. An intuitive verification of such background value is shown in the Fig. 2 (Left), where the “Label Token” curve has a high value in layer 0 with $\epsilon(p)=0$. However, the background value also indicates that the representation generated by BGE correctly clusters the samples, confirming its reliability.
@@ A.3 Training and Inference of Centroid Classifier
In this paper, we follow Cho et al. (2024) to train centroid classifiers as a probe toward hidden states of LMs. In detail, given the LM’s hidden states set $\left\{h_{i}^{l}\right\}_{i=1}^{m}$ of the selected tokens (according to the experimental setting, the last label token or forerunner token) in layer $l$ from an [ICL input]-[query label] set $\mathcal{Z}=\left\{\left(p_{i},y_{i}\right)\right\}_{i=1}^{m}$, where the labels are limited in label space $\mathbb{Y}$, in the training phase, we calculate the centroid of the hidden state $\bar{h}^{l}_{y}$ for each label respectively:
@@
Implementation. In our experiments, we set training sample number $m\fallingdotseq 256$, similarity function $\left\langle a,b\right\rangle\fallingdotseq-\|a-b\|_{2}$.
@@
In Fig. 6 (Middle), we define a Correct Label Assignment, here we introduce how this measurement is calculated. Suppose we have an attention score $\mathcal{A}_{W,f}(K,Q)$ calculated as (also written as $\alpha_{K\rightarrow Q}$ in brief):
@@
Figure 10: Visualization of non-trivial attention connection defined and disconnected in ablation analysis in §5.1 and Table 1. Notations and line numbers are same as Table 1.
@@
For Fig. 8, we calculate the magnitude of Step 1 as the finite differences of kernel alignment in Fig. 2 (Left, “Forerunner Token of Label”). We directly use the head counting of Fig. 5 (Middle) and Fig. 6 (Left) as the magnitude of Steps 2 and 3. These data are regularized and converted into transparencies.
@@ A.7 Experiment Setting of Ablation Experiments in Table 1
For reference, for each experiment, we also conduct controlled experiments where the same amount of randomly selected attention connections are removed from the same layers as the controlled values. The controlled results are shown as smaller numbers under the experimental results.
@@ Appendix B LM Pre-training Dynamics Measured by ICL Circuit
Figure 11: Operating magnitude for each inference step (normalized) and ICL accuracy against pre-training steps on Pythia 6.9B and SST-2.
One can divide a self-regression model into an early part and a later part in depth of layers, where the early part encodes the input into a hidden representation, and the later part decodes the hidden representation back to the input. So, the training object can also be divided into an encoding loss and a decoding loss. According to the discussion in §6, the operation of Step 1 can be classified as encoding, and the other two steps of the induction circuit can be classified as decoding. Intuitively, since the decoding operations require the encoding results as input, unless the encoding operation converges to a stable output, the decoding can not be trained since the input-output mapping is noised, causing unstable gradients to interfere with the training (Liu et al., 2020).
We confirm such a hypothesis by measurements in Pythia 6.9B (Biderman et al., 2023) as shown in Fig 11, where: (1) The magnitude of the 3 operations emerge in the early phase of pre-training (less than 10k steps), and is monotonically increasing, while the encoding operation has the fastest growth rate. (2) ICL capacity appearance after all three operations reaches a high level (around 50k steps, notice that the random accuracy is 0.5), while the curve morphologies of the operating magnitudes against the layer numbers (shown in Fig. 12) are convergence to the last training step. Such results suggest that: LMs start to produce the inner encoding in the very early steps of the pre-training, and can be an important fundamental in building the subsequent induction circuits, as explained in the previous works mentioned in §2.1.